Simplify log (base n) n^812
Answer by jessica43(140)

To solve this problem, you are going to need to use the base b logarithm formula to rewrite the problem:
log[base b]a = c means the same as b^c = a
So use this in your problem:
log[base n] n^812 = x
a = n^812, b = n, c = x
Now rewrite as an exponent using the log formula:
b^c = a
n^x = n^812
x = 812
So log[base n] n^812 = 812